If the common difference of an AP is 3, then what is a15 - a9 ?

Grade:10

2 Answers

Yash Baheti IIT Roorkee
9 years ago
Hi,

If d=3 than A15 will be = a + 14d (a being first term)
Similarly A9 will be + a + 8d,

S0 A15-A9 will be = 6d OR 18 (as d is 3)
Shivam
6 years ago
First of all if d=3 and we have to solve this then a15=a+14dAnd a9 =a+8dWe put this equation (a+14d)-(a+8d)=14d-8d=6d(6×3=12)